Hi I really need help here!
The sum of Elisa's ageand Ira's age is 47. Ira's age is 5 years less than three times Elia's age. How old are Ira and Elisa?

Answer by stanbon(75887)

The sum of Elisa's ageand Ira's age is 47. Ira's age is 5 years less than three times Elia's age. How old are Ira and Elisa?
---------------
EQUATIONS:
E + I = 47
I = 3E -5
-------------
Rearrange:
E +I = 47
3E -I = 5
---------
solve by any method to get:
E = 13 (Elisa's age)
I = 34 (Ira's age)
=================
Cheers,
Stan H.